2. For the list a - 1-1,8,21,tuple b-(5,1,-3), and ndarray c -np.array (11,1,21) determine the results of the following expre Python please
Answer #1

Python 3.7.0 Shell File Edit Shell Debug Options Window Help after a + [-2] list looks like this [4, 7.0, 4.0, -2] after a +Fle Edit Format Run Options Window Help for i in range(len(a)): alil ali]+3 printf a +3-) printfafter a[o] 4 list looks likProblem 2py-CrUsers Dell Deaktop Problem 2py 13.70) File Edit Format Run Options Window Help a [-1,8,2] c- numpy.array(I1,1,2

import numpy
a = [-1,8,2]
b = (5,1,-3)
c = numpy.array([1,1,2])

for i in range(len(a)):
a[i] = a[i] * 0.5
print("a * 0.5 = ",a)

for i in range(len(a)):
a[i] = a[i] + 3
print("a + 3 = ",a)

a[0] = 4
print("after a[0] = 4 list looks like this",a)

a = a + [-2]
print("after a + [-2] list looks like this",a)

a = a + [-2]*2
print("after a + [-2]*2 list looks like this",a)

print("\n")

c = c * 0.5
print("c * 0.5 = ",c)

c = c + 3
print("c + 3 = ",c)

c[0] = 4
print("after c[0] = 4 list looks like this",c)

print("\n")

for i in range(len(b)):
b[i] = b[i] * 0.5
  
print("b * 0.5 = ",b)
you have any query then plz ask me without any hesitation in the comment section below, if you like my answer then please thumbs up for the answer, before giving thumbs down please discuss the question it may possible that we may understand the question different way and we can edit and change the answers if you agree, thanks :)